OpenStudy (anonymous):

5 3/4 + 5 3/4 = please break down

OpenStudy (anonymous):

\[5\frac{3}{4} + 5\frac{3}{4}\]?

OpenStudy (anonymous):

\[= \frac{46}{4}\]

OpenStudy (aroub):

=23/4 +23/4 =46/4 =23/2

OpenStudy (anonymous):

5 3/4 + 5 3/4 please help break down answer to understand better

OpenStudy (anonymous):

yes